What is the chance of drawing a red marble when there are 12 marbles and three are red?
The chance is 25%.
How many outcomes are there if you roll two six- sided number cubes?
There are 36 possible outcomes. But if the cubes are identical, then for every
possible outcome, there's another one that looks just like it, so only 18 that
you can identify.
What are the z score boundaries for the middle 60 percent of the distribution?
The boundaries are ± 0.8416
Is having a baby boy hurt more than having a baby girl?
No they all hurt. It's more about hos big the baby is and how long it takes, complications etc.
What is the z-score for a value of 125?
The z-score for an observation (or a set of them) can only be calculated if the mean and standard deviation are known. Neither of them are given in this question.
How many clubs are there in a standard card deck?
There should be 13 of each suit (Spades, Diamonds, Clubs, and Hearts). Ace through King in numerical order, where A=1, J=11, Q=12, and K=13.
There are 13 clubs in a deck of cards, there are 4 suits, each containing 13 cards.
And in a pack of 54, it includes all those plus two jokers.
What is the probability of choosing a higher card than the 6 of diamonds?
depends on your comparision method: do you compare ranks only, or also suits ?
It appears red because of how hot it is and how close it is to the sun.
Drew had 24 marbles more then Marcus
Marcus had 25-12= 13 to begin with
Drew had 25+12 = 37 to begin with
37-13 = 24 is the difference between Drew and Marcus
If we understand your statement, then we're forced to disagree with it.
It goes along fine until you reach the last few words. The alternatives approach
probabilistic ratios, but the ratios will never be predictable.
Perhaps they are, in principle, as the number of repetitious events considered
becomes larger. But we believe the hooker lies in that word "larger", because
probabilities only drive predictions when the number of trials is infinite.
So, "large" must be evaluated by comparison with "infinity", and when you do that,
you never get any closer, no matter how many events you consider.
What is the probility that the card you pick out is not a face card or an ace out of all of a deck?
There are 52 - 16 = 36 cards that are not what you described.
So the probability is 36 / 52
Is there a chance that Australia might cease to exist?
Certainly. There is a chance that the earth will cease to exist.
z = 1.52 (approx)
18 of them.
What are two exclusive foods in RI?
Rhode Island is known for its unique culinary offerings, particularly its coffee milk, a sweet syrup made from coffee and milk that's a local favorite. Another exclusive food is the Rhode Island-style clam chowder, which is clear broth-based and often includes quahogs, a type of clam native to the region, differing from the creamy New England clam chowder. Additionally, stuffies, or stuffed quahogs, are a popular dish featuring clams mixed with breadcrumbs and seasonings.
When a card is drawn from an ordinary deck WHAT IS the probability of getting a jack?
2 out of 52
this can be canceled down too 1 out of 26
also 26%
Actually is would be 4/52 since there are 4 jacks.
Not 26%...more like 1/26=.03846 or 3.8465 if the answer were 1/26 which it is not.... the answer is 4/52 or 1/13.
So...1/13=.076923 so 7.6923% actually.. :)